3) Spinelli Company is deciding whether to automate one phase of its production process. The equipment has a six year life and will cost $450,000. The interest rate is 12%. Net cash inflows per year:

Year 1

$85,000

Year 2

$70,000

Year 3

$95,000

Year 4

$75,000

Year 5

$85,000

Year 6

$96,000

a.What is the present value of the net inflow for year 1?

b.What is the present value of the net inflow for year 5?
Answer #1

a.What is the present value of the net inflow for year 1?
= $85000*1/(1+0.12)^1=85000*0.893
=$75905

b.What is the present value of the net inflow for year 5?,
=$85000*1/(1+0.12)^5=85000*0.567
=$48195